OpenStudy (anonymous):

Let f(x)=x3−2x and g(x)=2x3−2x+5. Which of the following is equal to f(x)−g(x)? −x3−4x−5 3x3−5 −x3−5 x3+5

OpenStudy (anonymous):

−x3−5

OpenStudy (zale101):

Subtract the function f(x) and g(x) f(x)=x3−2x and g(x)=2x3−2x+5. \((x3−2x) -(2x3−2x+5)=?\)

OpenStudy (zale101):

@leon549 Hello, it would be great if you give explanation to your result, that would help. However, avoid giving direct answers. Thank you.

OpenStudy (anonymous):

I know right.. thank you.

OpenStudy (anonymous):

sry